MICQ(7) MICQ(7) mICQmicq-.on-line,. MICQ verbose[] mICQ-,.(): 8. 16. 32. 64. 256v6. 4096v7/v8. 8192v7/v8. 16284. 65536--. 131072 --. 262144 . 2097152 a--. 4194304 /. ,0,1. clear. sound[on|off|event] on()off(),event()"".0.4.10. autoaway[|on|off] "",on(),off().0off().,. auto[on|off] .on()off(). auto. alias[] .,.,%s,,,.0.4.10. unalias .0.4.10. alter.aliasmICQ-0.4.10. trans[|]... ,iso,.LANG,LCALLLCMESSAGES,!,auto()default()..,none() unload().-,debug."",@fun. ˜ /.micq/i18n/*.i18n,"". ,,.177,"transde117default". mICQ 1 MICQ(7) MICQ(7) uptime mICQ. seton()off(),,.: hermit: -, delbs:deletebackspace-, funny:

"",

log:, logonoff: , auto:,."", uinprompt: , autosave: micqrc, autofinger: UIN-, linebreak: simple,():,break,():,indent,():,smart,():; tabs:simple,():,cycle,():,,cycleall,():(),; silent::off(),on(),,complete(),,. 0.4.10. save,˜ /.micq/micqrc,.,. q[uit] exitmICQ-.quitexit0.4.10. <>,.UIN-.,. msg<>[] ..,(.).,""(#). .<<<<<<>>>>>>;,===.<<<<>>>>icq8(-2);,--=.,<<<>>>icq5, -1,-4.. ;,.,,:-2..,,.,-4. (UTF-8),,.,,,,().-1,-4,. a[].msg. r[].msg. url<>url url-. sms[][] SMS,+<><>,,..--. getauto[auto|away|na|dnd|occ|ffc][] ,away"",na"",dnd"",occ""ffc"".auto(),..0.4.10. auth[req|grant|deny|add]<> ,-,-. mICQ 2 MICQ(7) MICQ(7) resend<> . last[<>] ,. tabs.simple().tabsset. rand. s[<>] ,. e,w,ee,ww,eg,wg,eeg,wwg -."": +.ww. #,UIN.www. *. -. ˆ :. wwee"".,ˆ . &. |. :. ˆ ,IP. e-. w-. ee-,. ww-,.. eg-,-.0.4.10. wg-,-.0.4.10. eeg-,,-.0.4.10. wwg-,.,-.0.4.10. ewide-,. wide-,. status[] .IP,ICQ,UIN-a,,. :;sww. f[inger]uin| infouin| uin|. i. search mICQ 3 MICQ(7) MICQ(7) search@ search search @,(@),.,,,,,. add[group][] -,addgroup.0.4.10. :save(""). add[alias] add[alias]uin uin--,.(addalias)0.4.10. :save(""). rem[group][all] rem[group] -,-all().0.4.10.:save(""). rem[alias][all] .,all(),. (remalias)0.4.10.:save(""). togig. toginv . togvis contacts. ICQ reg. passicq. :o´ (0xf3). :save("")˜ /.micq/micqrc(micqrc(7)),. change[[]] .,.. online"". away[]

""..

na[]"".. occ[]"".. dnd[]"".. ffc[]"".. inv"". update (,.). other. about. setr[].,. mICQ 4 MICQ(7) MICQ(7) meta[show|load|save|set|get|rget]<> -.: show-. load-. save-. set-. get-. getr-,. 0.4.10. file[...] peerfile.0.4.10. peeruin| UIN-uin. open--TCP. close--. off. file. files[]... .,.’/’,,’.’. conn[] ,. open. login. close.. remove (). select.UIN. contact[] -: show-;. diff-(uin/). import -. . peeknick.()ICQ;"""". micq(1),micqrc(5) (JamesMorrison)<ja2morrison@student.math.uwaterloo.ca>mICQ.: <lecicct@ptt.yu>. mICQ 5